Evaluate the line integral βˆ«π’–(π‘₯, 𝑦, 𝑧) Γ— ⅆ𝒓 𝐢 , where 𝒖(π‘₯, 𝑦, 𝑧) = (𝑦 2 , π‘₯, 𝑧) and the curve π‘ͺ is described by 𝒛 = 𝑦 = 𝑒 π‘₯ with π‘₯ ∈ [0,1].


1
Expert's answer
2021-12-15T10:14:48-0500

parameter for of C,

x=t, y=etΒ ,z=etΒ .

r(t)=t i+ etΒ j+etΒ k

=> dr=(i + etΒ j + etΒ k)dt

Then,

"\int_0^1 u.dr\\
=\int_0^1 (e^{2t},t,e^{t}).(1,e^t,e^t)dt\\
=\int_0^1 (2e^{2t}+te^{t})dt\\
=[ e^{2t}+te^{t}-e^t]_0^1\\
=e^2+e-e-1-0+1\\
=e^2"